A random sample of size 144 is drawn from a population whose distribution, mean, and standard deviation are all unknown. The summary statistics are š‘„Ģ…= 58.2 and s = 2.6.

a. Construct an 80% confidence interval for the population mean Ī¼.

b. Construct a 90% confidence interval for the population mean Ī¼.


1
Expert's answer
2022-06-02T13:58:27-0400

a. The critical value forĀ "\alpha = 0.20, df=n-1=143"Ā degrees of freedom isĀ "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=1.2875"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(58.2-1.2875\times\dfrac{2.6}{\sqrt{144}}, 58.2+1.2875\times\dfrac{2.6}{\sqrt{144}})"


"=(57.921, 58.479)"

Therefore, based on the data provided, theĀ 80%Ā confidence interval for the population mean isĀ "57.291 < \mu < 58.479,"Ā which indicates that we areĀ 80%Ā confident that the true population meanĀ "\mu"Ā is contained by the intervalĀ "(57.921, 58.479)."


b. The critical value forĀ "\alpha = 0.10, df=n-1=143"Ā degrees of freedom isĀ "t_c\u200b=z_{1\u2212\u03b1\/2;n\u22121}=1.655579"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})""=(58.2-1.655579\times\dfrac{2.6}{\sqrt{144}},""58.2+1.655579\times\dfrac{2.6}{\sqrt{144}})"



"=(57.841, 58.559)"

Therefore, based on the data provided, theĀ 90%Ā confidence interval for the population mean isĀ "57.841 < \mu < 58.559,"Ā which indicates that we areĀ 90%Ā confident that the true population meanĀ "\mu"Ā is contained by the intervalĀ "(57.841, 58.559)."